sophiaespinosa:

A ball has a mass of 140 g. What is the force needed to accelerate the ball at 25 m/s2? (Formula: F=ma) A.) 3.5 N B.) 115 N C.) 165 N D.) 4.5 N

Shadow:

\[F = ma\]

Shadow:

You have mass and acceleration, input them and calculate the force.

sophiaespinosa:

mmmm ok

sophiaespinosa:

never mind figured out the answer 3.5

Shadow:

k I'll just write out the explanation in case anyone else comes across this question

Shadow:

Before using the formula F = ma, first you need to convert your units Whilst the mass is given, at 140 grams, it is not in the unit of mass that F = ma. It wants mass in kg. So we convert. \[\frac{ 140g }{ 1 } \times \frac{ 1 kg }{ 1000 g } = 0.14kg\]

Shadow:

Then we do, \[F =ma \rightarrow F = 0.14 \times 25 = 3.5 N\]

sophiaespinosa:

ok thanks